OpenStudy (anonymous):

The perimeter of a face of a cube is 16 mm. What is its volume

OpenStudy (anonymous):

Since the perimeter is 16 mm, the length of the side of the cube is 16/4 = 4 mm. Then the volume is the length of the side cubed, so we get: 4^3 = 64 mm^3

OpenStudy (anonymous):

let side of the square =a For a square perimeter=4a area=a^2 volume=a^3